# Enhancing LLM Reasoning with Iterative DPO: A Comprehensive Empirical Investigation

### v2

@@ Appendix E More Results of DPO-VP Analysis
Furthermore, we conduct evaluations on two additional Out-of-Domain datasets: CMATH, a Chinese mathematics dataset, and MMLU_STEM, which covers non-mathematical subjects such as physics, chemistry, and biology. The results are presented in Table 8. Eurus-2-7B-PRIME, despite being fine-tuned on a large number of mathematical problems using only English training data, exhibits significant improvements in Chinese reasoning capabilities. However, this improvement comes at the cost of weakened performance in question answering across other domains. Meanwhile, other RL methods, as well as our Qwen2.5-7B-DPO-VP, which leverage a self-augmentation fine-tuning paradigm, demonstrate overall enhancements in both Chinese responses and non-mathematical domain questions.
pass@1 acc CMATH MMLU_STEM
Qwen2.5-Math-7B 64.0 58.3
Eurus-2-7B-PRIME 72.5 42.8
Qwen2.5-7B-Simple-RL 70.3 59.1
Qwen2.5-7B-PURE-VR 69.5 58.8
Qwen2.5-7B-DPO-VP 68.5 59.3
Table 8: Performance comparison on Out-of-Domain datasets.
@@ F.2 Iterative DPO Can Be Viewed as a High Off-Policy Degree Online RL
During the period of conducting this research, we observed that leveraging iterative DPO for model optimization has already become a viable low-cost alternative to RL fine-tuning for some enterprises and research institutions when extreme performance is not the primary goal. We are delighted by this coincidence, as it indirectly validates that we have caught up with the research frontier in the field of LLM reasoning. We also extend our gratitude to Zhang et al. (2025a), Liang Wen (2025), and all open-source contributors in the LLM community for their selfless contributions to the open-source ecosystem!

### v3

@@ Appendix E More Results of DPO-VP Analysis
E.1 Token Length and Self-Reflection Vocabulary Analysis
@@
E.2 Comparison with Long-Chain SFT Methods
Recently, reinforcement learning methods built upon long-chain SFT have received increasing attention. We compare our approach with two recent methods that perform long-chain SFT using limited data: LIMO (Ye et al., 2025) and S1 (Muennighoff et al., 2025). Since both released models are only available at the 32B scale, we reproduce their settings by applying SFT on our base model, Qwen-2.5-Math-7B, using their provided datasets. We follow their default hyperparameters and train for 15 epochs. During evaluation, we sample with $t=0.6$ and set the maximum response length to 32K tokens (in contrast, DPO-VP uses 2K tokens). The performance on standard benchmarks is summarized in Table 9.
Surprisingly, DPO-VP outperforms these limited-data SFT methods on nearly all evaluated benchmarks. One possible explanation is that, at the 7B scale, SFT with small-scale long-chain data may offer less performance gain compared to its effect on larger models like 32B. Nevertheless, both LIMO and S1 still yield moderate improvements over the base model.
pass@1 acc MATH500 Minerva Olympaid AMC23 AIME24 Average
Qwen2.5-Math-7B 64.8 15.4 25.6 37.5 16.7 32.0
Qwen2.5-Math-7B-LIMO 70.2 21.0 37.0 45.0 16.7 38.0  (+6.0)
Qwen2.5-Math-7B-S1 72.0 32.4 37.5 55.0 13.3 42.0 (+10.0)
Qwen2.5-7B-DPO-VP 74.8 35.3 36.9 67.5 26.7 48.2 (+16.2)
Table 9: Pass@1 Accuracy Results with Long-Chain SFT Methods.
E.3 Training Dynamics of DPO under Different Label Noise Levels
In addition to final performance, we also tracked the training dynamics of DPO under different noise levels. As shown in Figure 8, lower noise (e.g., $p\leq 0.3$) leads to higher training accuracy and more stable reward separation between chosen and rejected samples. In contrast, high-noise setups (e.g., $p=0.7$ or $p=1.0$) cause convergence slow-down and reward collapse, confirming the importance of maintaining accurate preference signals during training.
Figure 8: Training dynamics of DPO under different label noise levels. Lower noise consistently leads to better optimization behavior and more stable reward separation.
E.4 Ablation Study on the DPO Sampling Temperature
In this section, we investigate the impact of sampling temperature on DPO performance. Our motivation stems from the observation that fixed low-temperature sampling may lead to insufficient exploration in later DPO iterations, resulting in early performance saturation. In our default setting, we used a constant temperature $t=0.7$ across all DPO rounds. To explore the potential of lightweight exploration, we introduced a simple temperature increase strategy: using $t=1.0$ for epochs 4–5 and $t=1.2$ for epochs 6–7. Table 10 reports the average accuracy across five benchmarks under both temperature settings.
Epoch Fixed $t=0.7$ Ours (temperature $\uparrow$)
3 44.5 44.5
4 44.6 44.7
5 44.3 46.5
6 46.2 48.2
7 45.5 47.6
Table 10: Comparison of average performance (pass@1 across five benchmarks) under fixed and increased sampling temperatures.
As shown, the increased temperature leads to noticeable performance gains starting from epoch 5, indicating improved sample diversity and preference pair quality. This suggests that even without using formal entropy-based exploration objectives, a simple sampling heuristic can effectively encourage diversity in preference selection and improve downstream learning. We believe this highlights the potential of lightweight exploration strategies in DPO, especially in scenarios where reward supervision is sparse or static.
E.5 Use of Distilled PRM Instead of a Fixed Verifier (e.g., DeepSeek-V3)
While large-scale verifiers such as DeepSeek-V3 (a 671B MoE model) provide high-quality supervision, they are computationally expensive to query. Each DPO iteration would require over 7.5K queries per epoch if DeepSeek-V3 were used directly, which is prohibitive for multi-round training. To address this, we distill its supervision into a smaller PRM initialized from Qwen2.5-7B, enabling scalable and efficient learning. Table 11 compares three one-shot filtering variants, showing that PRM achieves nearly identical performance to DeepSeek-V3 filtering.
Setting GSM8K MATH500 Gaokao Minerva Avg
+Outcome Label 89.8 74.2 62.9 25.0 63.0
+Qwen2.5-PRM 90.5 74.6 61.0 25.0 62.8
+DeepSeek-V3 90.3 75.1 61.7 25.6 63.2
Table 11: Performance of one-shot filtering with different verifiers.
This result supports the use of a lightweight, distilled verifier. Additionally, our formulation generalizes to settings where a fixed verifier is available—e.g., we use outcome labels as a special case of verifier supervision in Section 3.5.
E.6 Comparison to Process-Level RL Methods
We include additional comparisons to strong RL baselines that use learned reward models, such as PURE-PRM, a PPO-based method trained with PRM800K. Results in Table 12 show that our method achieves comparable or better performance with less training complexity. We also report results for Eurus-2-7BPRIME and rStar-Math-7B in Table 2, both of which adopt full PRM-based supervision. Their performance does not consistently surpass ours.
Method GSM8K MATH500 Gaokao Minerva Avg
Ours (3 epochs) 91.5 75.0 65.8 34.2 66.6
PURE-PRM 88.3 79.8 63.9 34.1 66.5
Table 12: Comparison with PURE-PRM (process-level PPO with PRM800K).
E.7 About Using Monte Carlo Rollouts in PRM Training
Rather than training PRMs via Monte Carlo (MC) rollouts, which are costly and require full reward computation, we directly label sampled responses online using a strong verifier. This reduces computation and aligns well with our iterative, feedback-driven optimization.
To verify this design, we compare our online PRM with a variant trained on the public Math-Shepherd dataset. On ProcessBench in Table 5, our PRM achieves higher average F1 (73.6 vs. 67.2), suggesting that strong online supervision offers both quality and efficiency.
This strategy is also aligned with recent work (Chen et al., 2025), which uses bi-directional verifier signals to label process-level trajectories. While MC-based training remains more theoretically rigorous, our framework focuses on co-evolving generator and verifier under verifiable supervision. Therefore, this practical approximation is sufficient to support the findings of this work.
E.8 Evolution of Reasoning in DPO-Optimized Models
To better illustrate the behavioral differences between models, we present a case study comparing four model variants on a mathematical reasoning task involving the apparent magnitude of a star cluster (Figure 9). The task requires combining knowledge of logarithmic scaling and magnitude-distance relationships, and serves as a strong testbed for verifying the model’s ability to compute, reason, and self-correct.
@@ Appendix F Discussion: Why Iterative DPO Leads to Improvement
Based on the rigorous theoretical analysis in Xiong et al. (2024), intuitively, iterative DPO leads to improvement because it allows the model to continually refine its policy through updated feedback, thereby reducing reward misalignment over time. Additionally, the batch exploration strategy promotes diverse sampling, which enhances generalization and mitigates overfitting to early-stage preferences. The performance gains of iterative DPO on math tasks stem from the inherent self-reflection ability in the base model and the iterative nature of the near-online RL setting
@@ F.2 Iterative DPO Can Be Viewed as a High Off-Policy Degree Online RL
During the period of conducting this research, we observed that leveraging iterative DPO for model optimization has already become a viable low-cost alternative to RL fine-tuning for some enterprises and research institutions when extreme performance is not the primary goal. We are delighted by this coincidence, as it indirectly validates that we have caught up with the research frontier in the field of LLM reasoning. We also extend our gratitude to Zhang et al. (2025a), Liang Wen (2025), and all open-source contributors in the LLM community for their selfless contributions to the open-source ecosystem.